# Estimating Privacy Leakage of Augmented Contextual Knowledge in Language Models

### v2

@@ 4.4 Token $n$-gram Influence of Context
In this section, we analyze the context influence of each $i$-th token $n$-gram in the context $\hat{\tau}_{i,n}(\overline{p}_{\theta,\lambda})$, i.e., we compare the output probability with and without the $i$-th token $n$-gram from the context to measure the influence. Due to the possibly large number of token $n$-grams, we only evaluate 100 contexts. Figure 3 shows the results for various token (128, 32, 8, 4)-gram influence on PubMedQA for OPT-1.3B with $\lambda=1.0$. We observe two trends: (1) Larger $n$-grams have higher peak context influence, which is intuitive given that the more information (larger $n$) is removed from the context, the more likely the output of the LM will change; (2) for seemingly all $n$, the context influence peaks for earlier $i$-th token $n$-grams, i.e. small $i$, then gradually declines for later $i$-th token $n$-grams. The results suggest that the model is influenced by information located earlier in the context than those located late, which might stem from the larger issue of position bias Liu et al. (2024). This implies that practitioners who want to control the influence of certain sequences can place privacy-sensitive ones toward the end of the context.
@@ 5 Related Works
Parametric Knowledge Leakage. It has been demonstrated that inadvertent memorization of pre-training data can lead to privacy leakage Carlini et al. (2019); Song and Shmatikov (2019) in the form of extraction attacks Carlini et al. (2021); Thomas et al. (2020). Hence, there is extensive research on understanding the memorization dynamics of LMs Tirumala et al. (2022); Zhang et al. (2023); Lesci et al. (2024); Biderman et al. (2024), where it has been shown that various factors such as model size, data duplication, and prompt length increase memorization. From these results, works have proposed dataset curation techniques, such as data deduplication Kandpal et al. (2022), to mitigate training data privacy leakage. In this work, we seek to conduct a similar analysis for augmented contexts.
Contextual Knowledge Leakage. Recent works have demonstrated that LMs can leak privacy-sensitive information provided to a prompt during inference via prompt regurgitation Wang et al. (2023); Priyanshu et al. (2023). In particular, recent works have shown through the lens of contextual integrity theory Nissenbaum (2009) that LMs lack the ability to effectively reason about the information sensitivity of contextual knowledge Mireshghallah et al. (2024); Bagdasarian et al. (2024); Shao et al. (2024). On the other hand, our analysis operationalizes exp-post per-instance DP to understand the factors that unintentionally influence contextual knowledge leakage. Zeng et al. (2024); Qi et al. (2024) investigated attacks on RAG systems that extract contextual knowledge, a similar setup and goal to our work. However, these works focus primarily on data extraction and implicitly assume that the retrieved contexts from the RAG database are not contained in the LM’s parametric knowledge, which overly-attributes the privacy leakage to the contexts. Also related, Huang et al. (2023) investigated the privacy leakage of retrieval-based LMs, such as kNNs. Lastly, another body of work investigated MIAs for augmented contexts Anderson et al. (2024); Wang et al. (2024); Li et al. (2025). Context influence can be viewed as inferring membership of a context.
Context hallucination. Our work follows prior work on summarization factuality where the response from an LM conflicts with an augmented context Maynez et al. (2020); Pagnoni et al. (2021). We focus on hallucination mitigation during inference by utilizing PMI to amplify focus on contextual rather than parametric knowledge Van der Poel et al. (2022); Shi et al. (2023). Our results demonstrate how these decoding methods affect the privacy leakage of contextual knowledge. Another body of work Fernandes et al. (2021); Sarti et al. (2023); Cohen-Wang et al. (2024); Du et al. (2024) measured an LM’s reliance on an augmented context; however, the goal of these works is not motivated by privacy, and hence, their results/discussion are orthogonal to ours.
@@ Appendix B Proof of CID satisfying $\epsilon$-DP
We will now show how CID can satisfy $n$-gram level $\epsilon$-DP (Definition 2.1). First, we are going to slightly modify CID by first selecting $\lambda$ so that we bound the amount of information leaked from a context $D$ when releasing the next token $y_{t}$. The algorithm can be found in Algorithm 1, which follows from Husain et al. (2020); Flemings et al. (2024).
@@ Appendix C Additional Experimental Setup
{mdframed}
[backgroundcolor=gray!5, linecolor=black, linewidth=1pt]
PubMedQA
Document: Programmed cell death (PCD) is the regulated death of cells within an organism. The lace plant (Aponogeton madagascariensis) produces perforations in its leaves through …
@@
{mdframed}
[backgroundcolor=gray!5, linecolor=black, linewidth=1pt]
CNN
News article: (CNN)The Palestinian Authority officially became the 123rd member of the International Criminal Court on Wednesday, a step that gives the court jurisdiction over alleged crimes …
@@
{mdframed}
[backgroundcolor=gray!5, linecolor=black, linewidth=1pt]
PubMedQA
Document: .
@@
{mdframed}
[backgroundcolor=gray!5, linecolor=black, linewidth=1pt]
CNN
News article: .
@@
For hardware, all of our experiments used one A100 40GB GPU, except for the experiments using 30B and 66B which used two and three A100 40GB GPUs, respectively. Each experiment usually takes around 15 mins to run for OPT-1.3B, except for the $i$-th token $n$-gram experiments (Section 4.4 which take longer depending on how small the token $n$-gram is. For software, our summarization quality evaluation is based on the code from Xu (2023), which is freely available on GitHub  [^2] https://github.com/zhichaoxu-shufe/context-aware-decoding-qfs. All datasets and models used in our experiments are freely available at Hugging Face, and our research does not conflict with their intended use cases, which is to evaluate text generation quality and privacy. The CNN-DM dataset follows the apache-2.0 license, LLaMA 3 follows the Llama 3.2 Community License Agreement, which we agreed to before evaluating, and GPT-Neo follows the MIT license, all of which we ensured not to go against.
@@ Appendix D Additional Experimental Results
000.250.250.250.250.50.50.50.50.750.750.750.75111170707070110110110110150150150150190190190190temperature $T$$\mathbb{E}[f_{\text{infl}}(\overline{p}_{\theta})]$
000.250.250.250.250.50.50.50.50.750.750.750.75111115151515171717171919191921212121temperature $T$ROUGE-L
000.250.250.250.250.50.50.50.50.750.750.750.75111125252525323232323939393946464646temperature $T$FactKB
@@
$\lambda=1.5$ Japanese can sleep soundly in their beds tonight as the government’s top military official earnestly revealed that the country’s Air Self Defense Force (ASDF) had never encountered an extraterrestrial unidentified flying object.
$\lambda=1$ in a interview with Japanese defense minister, politician Antonio Inoki asked the defense minister about aliens and UFOs and the defense minister answered that the Air Self Defense Force (ASDF) has never encountered one.

### v3

@@ 4.4 Token $n$-gram Influence of Context
In this section, we analyze the context influence of each $i$-th token $n$-gram in the context $\hat{\tau}_{i,n}(\overline{p}_{\theta,\lambda})$, i.e., we compare the output probability with and without the $i$-th token $n$-gram from the context to measure the influence. Due to the possibly large number of token $n$-grams, we only evaluate 100 contexts. Figure 3 shows the results for various token (128, 32, 8, 4)-gram influence on PubMedQA for OPT-1.3B with $\lambda=1.0$. We observe two trends: (1) Larger $n$-grams have higher peak context influence, which is intuitive given that the more information (larger $n$) is removed from the context, the more likely the output of the LM will change; (2) for seemingly all $n$, the context influence peaks for earlier $i$-th token $n$-grams, i.e. small $i$, then gradually declines for later $i$-th token $n$-grams. The results suggest that the model is influenced by information located earlier in the context than those located late, which might stem from the larger issue of position bias liu2024lost. This implies that practitioners who want to control the influence of certain sequences can place privacy-sensitive ones toward the end of the context.
@@ 5 Related Works
Parametric Knowledge Leakage. It has been demonstrated that inadvertent memorization of pre-training data can lead to privacy leakage carlini2019secret; song2019auditing in the form of extraction attacks carlini2021extracting; thomas2020investigating. Hence, there is extensive research on understanding the memorization dynamics of LMs tirumala2022memorization; zhang2023counterfactual; lesci2024causal; biderman2024emergent, where it has been shown that various factors such as model size, data duplication, and prompt length increase memorization. From these results, works have proposed dataset curation techniques, such as data deduplication kandpal2022deduplicating, to mitigate training data privacy leakage. In this work, we seek to conduct a similar analysis for augmented contexts.
Contextual Knowledge Leakage. Recent works have demonstrated that LMs can leak privacy-sensitive information provided to a prompt during inference via prompt regurgitation wang2023decodingtrust; priyanshu2023chatbots. In particular, recent works have shown through the lens of contextual integrity theory nissenbaum2009privacy that LMs lack the ability to effectively reason about the information sensitivity of contextual knowledge mireshghallah2024can; bagdasarian2024airgapagent; shao2024privacylens. On the other hand, our analysis operationalizes exp-post per-instance DP to understand the factors that unintentionally influence contextual knowledge leakage. zeng2024good; qi2024follow investigated attacks on RAG systems that extract contextual knowledge, a similar setup and goal to our work. However, these works focus primarily on data extraction and implicitly assume that the retrieved contexts from the RAG database are not contained in the LM’s parametric knowledge, which overly-attributes the privacy leakage to the contexts. Also related, huang2023privacy investigated the privacy leakage of retrieval-based LMs, such as kNNs. Lastly, another body of work investigated MIAs for augmented contexts anderson2024my; wang2024membership; li2025generating. Context influence can be viewed as inferring membership of a context.
Context hallucination. Our work follows prior work on summarization factuality where the response from an LM conflicts with an augmented context maynez2020faithfulness; pagnoni2021understanding. We focus on hallucination mitigation during inference by utilizing PMI to amplify focus on contextual rather than parametric knowledge van2022mutual; shi2023trusting. Our results demonstrate how these decoding methods affect the privacy leakage of contextual knowledge. Another body of work fernandes2021does; sarti2023quantifying; cohen2024contextcite; du2024context measured an LM’s reliance on an augmented context; however, the goal of these works is not motivated by privacy, and hence, their results/discussion are orthogonal to ours.
@@ Appendix B Proof of CID satisfying $\epsilon$-DP
We will now show how CID can satisfy $n$-gram level $\epsilon$-DP (Definition 2.1). First, we are going to slightly modify CID by first selecting $\lambda$ so that we bound the amount of information leaked from a context $D$ when releasing the next token $y_{t}$. The algorithm can be found in Algorithm 1, which follows from husain2020local; flemings2024differentially.
@@ Appendix C Additional Experimental Setup
PubMedQA Document: Programmed cell death (PCD) is the regulated death of cells within an organism. The lace plant (Aponogeton madagascariensis) produces perforations in its leaves through …
@@
CNN News article: (CNN)The Palestinian Authority officially became the 123rd member of the International Criminal Court on Wednesday, a step that gives the court jurisdiction over alleged crimes …
@@
PubMedQA Document: .
@@
CNN News article: .
@@
For hardware, all of our experiments used one A100 40GB GPU, except for the experiments using 30B and 66B which used two and three A100 40GB GPUs, respectively. Each experiment usually takes around 15 mins to run for OPT-1.3B, except for the $i$-th token $n$-gram experiments (Section 4.4 which take longer depending on how small the token $n$-gram is. For software, our summarization quality evaluation is based on the code from xu2023context, which is freely available on GitHub  [^2] https://github.com/zhichaoxu-shufe/context-aware-decoding-qfs. All datasets and models used in our experiments are freely available at Hugging Face, and our research does not conflict with their intended use cases, which is to evaluate text generation quality and privacy. The CNN-DM dataset follows the apache-2.0 license, LLaMA 3 follows the Llama 3.2 Community License Agreement, which we agreed to before evaluating, and GPT-Neo follows the MIT license, all of which we ensured not to go against.
@@ Appendix D Additional Experimental Results
[…]
00.250.250.50.50.750.75111515171719192121temperature TTROUGE-L
00.250.250.50.50.750.75112525323239394646temperature TTFactKB
@@
$\lambda=1.5$
Japanese can sleep soundly in their beds tonight as the government’s top military official earnestly revealed that the country’s Air Self Defense Force (ASDF) had never encountered an extraterrestrial unidentified flying object.
$\lambda=1$
in a interview with Japanese defense minister, politician Antonio Inoki asked the defense minister about aliens and UFOs and the defense minister answered that the Air Self Defense Force (ASDF) has never encountered one.